# How Much are LLMs Contaminated?

### v1

@@ 6 Conclusion
As LLMs rapidly evolve and the data used for training become more extensive, LLMs’ performances are inevitably biased by contamination. Therefore, in this survey, we investigated and structured the current landscape of contamination detection in LLMs. We introduced a new Python library named LLMSanitize which is a universal tool for quickly detecting contamination in various models and datasets. The extensive investigation and proposed tool could provide a universal foundation for future models when considering data or model contamination issues.
Appendix A Data Contamination Detection Results
Data contamination detection method ARC HellaSwag Winogrande GSM8K
Valid Test Valid Test Valid Test Test
GPT-2 style 10.70 11.26 13.88 13.97 0.24 0.62 11.30
GPT-3 style 0.67 0.51 0.62 0.92 0.00 0.11 2.73
Exact match 0.00 0.00 0.00 0.00 0.00 0.00 0.00
PaLM-style 0.33 0.43 0.00 0.00 0.00 0.00 0.00
GPT-4 style 0.67 0.43 0.19 0.38 0.00 0.11 0.30
Platypus-style 11.37 9.56 22.14 20.08 0.87 1.41 7.51
Table 4: Data contamination detection results for several baseline methods described in Section 2. We compare each dataset validation or test set against its training set, and report the percentage of data points flagged as contaminated. All results were obtained using the LLMSanitize library, without subsampling any dataset partition.
In Table 4, we benchmark several data contamination techniques on datasets from the OpenLLM Leaderboard following Tables 1, 3 and 2. We discard datasets which do not have a training set (MMLU and TruthfulQA), and perform data contamination analysis between the training set and the validation (if available) and test sets. We notice that ARC and HellaSwag show on average higher contamination with the training set, especially with GPT-2-style string matching and Platypus-style embeddings similarity.
Appendix B More Model Contamination Detection Results
LLM Average ARC HellaSwag MMLU TruthulQA Winogrande GSM8K
Llama-2-7B-base 0.566 0.001 0.847 0.690 0.913 0.945 0.000
Llama-2-7B 0.278 0.064 0.489 0.146 0.001 0.001 0.968
[…]
Qwen-1.5-7B 0.432 0.000 0.023 0.700 0.664 0.202 1.000
[…]
Mistral-7B 0.312 0.000 0.000 0.010 0.042 0.999 0.444
[…]
Gemma-7B 0.417 0.978 0.000 0.580 0.122 0.820 0.000
Table 5: P-value of the guided prompting (with ROUGE-L) contamination detection (Golchin & Surdeanu, 2023b) for four open-source LLMs in both their base and instruction-tuned versions on OpenLLM Leaderboard datasets. Shaded results are not statistically significant.
In Table 5, we reproduce the guided prompting contamination detection results from Table 1 while including the base version (immediately after pre-training) of each LLM. Some datasets such as ARC or TruthfulQA seem to have contaminated all pre-training datasets. We note that although they are consistent, contamination levels may change in both directions from pre-training to instruction-tuning: for instance Llama-2 becomes contaminated on TruthfulQA and Winogrande after instruction tuning ; while Gemma does not show contamination anymore after instruction tuning on MMLU or TruthffulQA. We highlight that designing contamination detection techniques robust to the changes in the LLM knowledge after undergoing instruction tuning and alignment remains an open research problem.

### v3

@@ 6 Conclusion
As LLMs rapidly evolve and the training data become more extensive, LLMs’ performances are inevitably biased by contamination. Therefore, in this survey, we investigated and structured the current landscape of contamination detection in LLMs. We introduced a new Python library named LLMSanitize which is a universal tool for quickly detecting contamination in various models and datasets. The extensive investigation and library could provide a universal foundation for future models when considering contamination issues.
Limitations
In this paper, we reviewed all work related to the detection of contamination. Contamination itself, as a general problem with LLMs, is larger than detection. For instance, our study scope does not focus on the mitigation of contamination, which we only briefly touch in Section 4.1. Mitigation is also of critical importance ; but we made the choice in this study to narrow down and focus on detection, which we believe makes our contribution more focused.
We included all works mentioning the issue of detecting or assessing contamination with language models from the last four years. Many of these works, for instance language models reports (Radford et al., 2019), are not focusing on the issue of contamination detection: these papers simply discuss this issue in a few paragraphs. Our survey also includes papers entirely dedicated to contamination detection, such as the LLM Decontaminator (Yang et al., 2023b). The technical depth of the work related to contamination detection across such differing types of works is very heterogeneous. This is not reflected in our survey as we chose to include all papers on the same level.
Appendix A LLMSanitize Usage
In this section we showcase how to use our Python library LLMSanitize. The library is recommended to be used with Python 3.9 and CUDA 11.8. The first step is to create the environment and install the library.
The following code should be run in a terminal:
1conda create --name llmsanitize python=3.9
2conda activate llmsanitize
3pip install llmsanitize==0.0.8
We first review how to use LLMSanitize for an open-data contamination detection setup. In this case, we compare two datasets and no LLM is involved. Let us take a simple setup: we compare HellaSwag as evaluation dataset against ARC as training dataset, sub-sampling 100 evaluation data points, and using GPT-2’s 8-grams collision as contamination detection method.
The following Python code could be run:
1import argparse
2from llmsanitize import OpenDataContaminationChecker
3
4parser = argparse.ArgumentParser()
5parser.add_argument("--train_data_name", type=str, default="allenai/ai2_arc")
6parser.add_argument("--eval_data_name", type=str, default="Rowan/hellaswag")
7parser.add_argument("--eval_set_key", type=str, default="validation")
8parser.add_argument("--text_key", type=str, default="ctx")
9parser.add_argument("--n_eval_data_points", type=int, default=100)
10parser.add_argument("--method", type=str, default="gpt-2")
11args = parser.parse_args()
12
13contamination_checker = OpenDataContaminationChecker(args)
14contamination_checker.run_contamination(args.method)
From there, the library will run the contamination detection and log results to a dedicated file. We refer the reader to all argument fields defined in the main.py file ; which among other properties, provide an option to stream over the (large) training dataset in order to limit memory usage for the use case of open-data contamination.
Next, we review how to conduct a closed-data contamination detection use case, involving an LLM and an evaluation dataset. For this setup, we apply the Guided Prompting method on Llama-2 on the ARC dataset. For this method (and all closed-data contamination detection methods except Sharded Likelihood), we make use of vLLM for efficient inference.
This requires us to first open a vLLM instance. Run the following shell script, which assumes access to 2 GPUs:
1export LD_LIBRARY_PATH=/usr/local/cuda/lib64:\$LD_LIBRARY_PATH
2export RAY_memory_monitor_refresh_ms=0;
3export CUDA_VISIBLE_DEVICES=0,1;
4server_type=vllm.entrypoints.openai.api_server
5
6python -m \$server_type \
7 --model <path_to_local_llama2> \
8 --tokenizer <path_to_local_llama2> \
9 --gpu-memory-utilization=0.9 \
10 --max-num-seqs=200 \
11 --disable-log-requests \
12 --host 127.0.0.1 \
13 --port <port_number> \
14 --tensor-parallel-size 2 \
15 --download-dir /export/home/cache
Next, a minimal Python code to run Guided Prompting detection could be:
1import argparse
2from llmsanitize import ClosedDataContaminationChecker
3
4parser = argparse.ArgumentParser()
5parser.add_argument("--eval_data_name", type=str, default="allenai/ai2_arc")
6parser.add_argument("--eval_set_key", type=str, default="test")
7parser.add_argument("--text_key", type=str, default="question")
8parser.add_argument("--label_key", type=str, default="anskwerKey")
9parser.add_argument("--n_eval_data_points", type=int, default=100)
10parser.add_argument("--num_proc", type=int, default=100)
11parser.add_argument("--method", type=str, default="guided-prompting")
12parser.add_argument("--local_port", type=int, default=<port_number>)
13parser.add_argument("--model_name", type=str, default=<path_to_local_llama2>)
14parser.add_argument("--guided_prompting_task_type", type=str, default="CLS")
15parser.add_argument("--user_local_model", type=bool, default=True)
16args = parser.parse_args()
17
18contamination_checker = ClosedDataContaminationChecker(args)
19contamination_checker.run_contamination(args.method)
Note that the model path and port number passed as arguments in the Python script should match the ones used to launch the vLLM instance.